OpenStudy (sugarapples):

find the geometric mean between each pair of numbers 11. 36,24 please help will fan them and give a medal

Directrix (directrix):

The geometric mean of 36 and 24 can be thought of as the square root of the product of the two.

Directrix (directrix):

If you are studying Geometry, I would think of it this way: 36 is to x as x is to 24 where x is the geometric mean. 36/x = x/24 x*2 = 36 * 23 x = ?

Directrix (directrix):

@sugarapples

Directrix (directrix):

What did you get for the answer?